Isamay0916:

The length of a rectangular frame is represented by the expression 2x + 8, and the width of the rectangular frame is represented by the expression 2x + 6. Write an equation to solve for the width of a rectangular frame that has a total area of 160 square inches. 4x2 + 32x + 48 = 0 4x2 + 28x − 112 = 0 2x2 + 32x − 112 = 0 x2 + 12x + 48 = 0

darkknight:

so do you know the formula for the area of a rectangle?

Isamay0916:

yes

Isamay0916:

A= wl

darkknight:

So area is length times width we are given that the length is 2x+8 and the width is 2x+6

darkknight:

So that means that 2x+8)(2x+6) = 160

darkknight:

Do you know how to factor?

Isamay0916:

yes i do

darkknight:

so to solve this, we will use foil. 2x+8)(2x+6

ramen:

yes so just first multiply the terms together foil

Isamay0916:

ok ty

darkknight:

which is 4x^2+16x+12x+48 which is 4x^2+28x+48

ramen:

this is not one of the solutions darkknight

ramen:

this is an example of a mistake that we can commonly make, so lets try and find the error he made

darkknight:

Are you sure the answer choices are right @Isamay0916

Isamay0916:

yes let me copy them down again.

ramen:

ok actually you may have maybe switched one of the numbers

Isamay0916:

The length of a rectangular frame is represented by the expression 2x + 8, and the width of the rectangular frame is represented by the expression 2x + 6. Write an equation to solve for the width of a rectangular frame that has a total area of 160 square inches. 4x2 + 32x + 48 = 0 4x2 + 28x − 112 = 0 2x2 + 32x − 112 = 0 x2 + 12x + 48 = 0

darkknight:

Did you copy the question right?

Isamay0916:

yes i did

ramen:

well the answer is definitely supposed to be 4x^2+28x+48 but it doesn't seem like the solutions are right

Isamay0916:

hmm thats strange

ramen:

yes it really is..

Isamay0916:

oh well thank you for helping me

Isamay0916:

@darkknight @ramen

darkknight:

no we are probebly doing something wrong

darkknight:

hold on

ramen:

i feel it too

Isamay0916:

ok

darkknight:

Got an idea So we have 4x^2+28x+48 but that is equal to 160 subtract 160 and you have 4x^2+28x-112

darkknight:

That is your answer

darkknight:

I am the darkknight

Isamay0916:

haha thank you darkknight!

ramen:

THATS WHAT I SAID DARKKNIGHT BUT YOU SAID I WAS WRONG

ramen:

mmmm lmao

Isamay0916:

lol

darkknight:

@ramen, where?

ramen:

aww shoot i deleted it bc i thought it was wrong

Isamay0916:

Was it this one "well the answer is definitely supposed to be 4x^2+28x+48 but it doesn't seem like the solutions are right"

darkknight:

okay ramen. I didn't see that answer

ramen:

yes because you werent paying attention but ok.

Isamay0916:

Okay, okay, both of you found the answer and ty

darkknight:

yw

ramen:

yw good job